a. Briefly explain two primary reasons for a firm to hold cash. (4 marks) b. A firms weekly average cash balances are as follows: Week 1 Week 2 Week 3 Week 4 RM24,000 RM34, 000 RM10,000 RM15,000 nual interest rate is 12 percent, how much return can be earned on the average (5 marks) cash balance? c. Using appropriate example, distinguish the lead and lag strategies. (6 marks) d Classity whether the following actions INCREASE, DECREASE, or CAUSE NO CHANGE to cash. This years tax liability is increased. Marketable securities are purchased with retained earnings. A sale is done based on the term 3/15, net 30 100,000 of new company shares floated in the open market. Raw material is purchased on credit (2 marks) (2 marks) (2 marks) (2 marks) (2 marks) iv. V. TOTAL 25 MARKS
Answer #1
a. Primary reasons to hold cash
1.To meet immediate day-to-day expenses & any sudden obligations
2.To meet any unforeseen need for hard-cash
b.
Week Balance/Addition Interest
1 24000 24000*12%*3/52= 166.15
2 10000 10000*12%*2/52= 46.15
3 -24000 -24000*12%*1/52= -55.38
4 5000
Bal.at end of Wk.4 15000 156.92
c. Lead & lag strategies are practised between subsidiaries, when give & take of cash involved.
That subsidiary which has excess cash,pays even before the due date for goods bought from the other subsidiary ,which has inadequate cash balance.This strategy is called LEADING.
Whereas,
when the former sells goods to the latter,it allows more time to the latter-sometimes, even to lag beyond due date--to make payment by relaxing its payment terms.This strategy is called LAGGING.
This helps in reducing the debt burden of the weaker subsidiary.
d.
i. This year's tax liability is increased-- No Change--- No cash outflow/Inflow
ii.Marketable securities are purchased with retained earnings--No Change--No cash outflow/Inflow
iii.A sale is done based on the term 3/15,net 30--- No change --it is a credit sale
iv.100000 of new co. share floated in the open market-- No Change-- cash is received only on subscription & purchase by the investor.
v.Raw material is purchased on credit--- No Change-- It is a credit or on account purchase.